# TreeLearn: A Comprehensive Deep Learning Method for Segmenting Individual Trees from Ground-Based LiDAR Forest Point Clouds

### v2

@@ Abstract
Laser-scanned point clouds of forests make it possible to extract valuable information for forest management. To consider single trees, a forest point cloud needs to be segmented into individual tree point clouds. Existing segmentation methods are usually based on hand-crafted algorithms, such as identifying trunks and growing trees from them, and face difficulties in dense forests with overlapping tree crowns. In this study, we propose TreeLearn, a deep learning-based approach for tree instance segmentation of forest point clouds. Unlike previous methods, TreeLearn is trained on already segmented point clouds in a data-driven manner, making it less reliant on predefined features and algorithms. Furthermore, TreeLearn is implemented as a fully automatic pipeline and does not rely on extensive hyperparameter tuning, which makes it easy to use. Additionally, we introduce a new manually segmented benchmark forest dataset containing 156 full trees, and 79 partial trees, that have been cleanly segmented by hand. The data is generated by mobile laser scanning and contributes to create a larger and more diverse data basis for model development and fine-grained instance segmentation evaluation. We trained TreeLearn on forest point clouds of $6665$ trees, labeled using the Lidar360 software. An evaluation on the benchmark dataset shows that TreeLearn performs equally well or better than the algorithm used to generate its training data. Furthermore, the method’s performance can be vastly improved by fine-tuning on the cleanly labeled benchmark dataset. The TreeLearn code is available from https://github.com/ecker-lab/TreeLearn. The data as well as trained models can be found at https://doi.org/10.25625/VPMPID.
@@
Correspondence: {jonathan.henrich, jan.vandelden}@uni-goettingen.de
@@ 1 Introduction
Today, forests are not only exposed to an ever-growing pressure from human use but also to a rapidly changing environment due to climate change. To understand changes in forest composition and structure as a response to these stressors, as well as for successful forest management, scientists and forest managers require precise and detailed information on the current status of our forests. To assess the structural characteristics of forests or individual trees, laser scanning techniques are a powerful tool [1, 2, 3]. They allow to create high-resolution 3D point clouds of the surfaces in a forest. To analyze such data, it is useful to separate the point cloud into the individual trees (Figure 1a), e.g. to classify them by their species [4, 5, 6] or to estimate the above-ground biomass [7, 8]. Separating a point cloud into individual trees is an instance segmentation problem: tree points must first be identified in the point cloud and then each point must be assigned to an individual tree instance. Since manually segmenting forest point clouds is highly time-consuming and subjective, there is a need for methods that automate this process. The method proposed in this work tackles this task for point clouds acquired using laser scanners located below the forest canopy. Primary sources of such point clouds are terrestrial and mobile laser scanning (TLS and MLS). Point clouds obtained by solely scanning from above the forest canopy usually do not capture the lower parts of the forest in high detail and therefore require a different set of methods, e.g. [9, 10].
@@
The automatic segmentation of MLS and TLS point clouds into individual tree instances is a relatively young research field that has so far been mostly tackled by algorithms based on hand-crafted features. These algorithms usually first identify tree trunks and then assign the remaining points to these trunks based on a fixed set of rules. In early works, these rules were relatively simple. For example, some authors divided the point cloud into clusters and merged them based on distance and relative orientation [11, 12]. Follow-up work building upon these methods included more elaborate local geometry and shape features [13]. More recently, [14] avoided complicated feature computations and proposed to cluster trees from bottom to top in a layer-by-layer manner. Another strand of literature conceptualized the forest as a graph and segmented trees via graph cut [15, 16, 17] or graph pathing methods [18]. Furthermore, some works aimed to incorporate biological theories. For example, [19] constructed features based on plant morphology theory which were then used in a region-growing approach. Similarly, [20] developed a shortest-path algorithm based on metabolic theory, which states that plants tend to minimize the transferring distance to the root. [21] also employed metabolic theory by constructing a superpoint graph where the node features depend on shortest path analysis. Recently, [22] proposed a segmentation method that identifies distinctive features in the forest and extracts individual trees based on topological structure. Compared to previous approaches, the method is more convenient for practitioners as it reduces the need for intensive hyperparameter tuning.
@@
Although modern frameworks achieved significant improvements in performance, automated tree instance segmentation remains prone to errors and usually requires manual postprocessing [23], especially in dense forests with heavily intersecting tree crowns. In such scenarios, hand-crafted features and pre-defined heuristics appear to be too inflexible to account for the plethora of possible interactions between trees. Machine learning approaches could overcome these limitations, since features and association rules are derived directly from the data through gradient-based learning. In fact, recent advances in point cloud processing have been dominated by machine learning methods, as indicated by their performance on various point cloud processing benchmarks, e.g. [24, 25, 26].
@@
Although machine learning methods have been applied to the segmentation of trees from forest point clouds  [27, 28, 29, 30], they mostly rely on projections into 2D. As a result, they lose information, because the network does not operate on the original point cloud. Other works do employ 3D deep learning for semantic segmentation, but resort to classical algorithms for tree instance segmentation [31, 32, 33, 34]. To our knowledge, there are only a few studies in the related field of urban tree segmentation that actually employ 3D deep learning methods for tree instance segmentation [35, 36]. Furthermore, one concurrent work [37] recently showed the potential of directly performing deep-learning-based instance segmentation on forest point clouds. It is surprising that this research direction has not been explored more thoroughly since state-of-the-art instance segmentation methods for 3D point clouds exclusively operate on 3D input data [38, 39, 40, 41, 42, 43], making them promising candidates for improving existing tree segmentation methods. In this paper, we adapt one of these methods [39] and propose TreeLearn, a comprehensive deep learning method to segment trees from forest point clouds. Our method first identifies tree points in the point cloud, then projects them towards the respective tree base, and groups them via density-based clustering. The whole process is realized in an automatic pipeline which can be used out-of-the-box without extensive hyperparameter tuning. This reduces the number of work steps required to be performed by the user to a minimum, which makes our method easy to apply.
@@
Since data-driven deep learning methods require large amounts of data for parameter optimization, high-quality labeled forest point clouds are urgently needed. Although there are some studies that use forest point clouds with manually corrected segmentation labels, some of them do not make their datasets public [16, 17]. Furthermore, those studies that provide publicly available datasets [44, 45, 46, 10, 47] often have properties which limit their use for training machine learning methods. For example, [44] only provides segmentation labels for a small number of trees located within a larger forest plot without segmentation labels. [45] acknowledge quality limitations of their tree segmentations as well as low laser scanning coverage. Other studies only provide labeled trees instead of the complete point clouds [46, 10]. We argue that machine learning methods should be able to handle the complete registered point clouds to make its usage as easy as possible for practitioners. Furthermore, the separation of tree points from understory and ground points is often non-trivial and therefore itself an interesting training objective. A concurrent work [47] acknowledges the need for complete labeled forest point clouds by providing a large-scale dataset with high-quality semantic and instance segmentations of 1130 trees. In our work, we also provide a benchmark dataset of a continuous forest with high-quality crown segmentations and classification of points into non-tree and tree. It contains 156 trees that are completely and 79 trees that are partly within the labeled forest area (Figure 1b). Since forest types as well as laser scanning characteristics are diverse, we argue that every labeled forest point cloud is an important step to create a rich data basis for the development of powerful and general machine-learning-based tree segmentation algorithms. Additionally, the dataset can be used to systematically evaluate and compare existing methods. In summary, our study has two main contributions:
@@
2 Dataset
In Section 2 and Section 3, we describe the dataset and our TreeLearn pipeline in detail. Readers unfamiliar with the technical background can skip directly to the results section.
2.1 Scanning Areas
The data basis for model training and validation consists of MLS point clouds from 19 forest plots located in Germany. Eight plots are located near the city of Göttingen, in Lower Saxony, two near Allstedt in Saxony-Anhalt, two near Oppershofen in Hesse, and seven near Lübeck in Schleswig-Holstein. The stands are all dominated by European beech (Fagus sylvatica L.) and between 92 and 162 years of age. A detailed description can be found in [48]. The size of the study plots ranged from 0.9 to 2.2 ha.
2.2 Data Acquisition
The plots were recorded in February 2021, in leaf-less condition, using a ZEB Horizon mobile laser scanner (Geoslam Ltd., Nottingham, UK). The device uses a laser to measure the distance and direction to surrounding objects while being carried through a scene. Based on the principle of simultaneous localization and mapping (SLAM) the ZEB Horizon creates a 3D point cloud of the scenery up to a distance of 100 m from the walking trajectory of the operator. Each hand-held walkthrough with the scanner resulted in a file that was used to create a point cloud by conducting the spatial coregistration (actual SLAM) in GeoSlam HUB Version 6 [49]. The resulting point clouds were subsampled to 1 cm resolution to reduce scan artifacts. Additionally, noise filtering was conducted to remove points that are not of interest, such as scanned particles in the air.
2.3 Segmentation Labels
To train the neural network in a data-driven manner, already segmented forest point clouds are required. We obtained our segmentation labels for the forest point clouds in two ways: (1) by using the automatic segmentation functionality provided by the Lidar360 software [50] and (2) by manually correcting the segmentation labels of a single forest plot. The automatic segmentation labels were used for training, while the manually corrected labels were used for fine-tuning and evaluation.
@@ 2.3 Segmentation Labels
To automatically segment the 19 forest point clouds into individual trees, we used the TLS-package from Lidar360 as described by [48]. The hyperparameters of the Lidar360 software were optimized to produce visually appealing tree segmentations. A minimum size of 10 m was chosen as a requirement for tree classification. Consequently, vegetation that is smaller than 10 m or points belonging to the forest ground were classified as non-tree points. Lidar360’s segmentation algorithm operates on a point cloud that is terrain-normalized, i. e. has a flat ground surface. This was achieved in a separate preprocessing step. To bypass the need for terrain normalization for our method, labels from the terrain-normalized data were transferred to the original data. The automatic segmentation contains numerous errors, including multiple trees detected as a single tree, non-tree points identified as part of trees, and inaccuracies in discerning tree boundaries in crowns and branches that are close to each other. Since these labels contain numerous mistakes, we refer to them as noisy labels. The noisy labels were used for training since individual trees are roughly captured correctly. The model obtained from training on these noisy labels serves as a starting point for fine-tuning.
Manually corrected segmentation
@@
For one forest plot, we manually corrected the segmentation labels generated by Lidar360 with the help of CloudCompare [51] by comparing each tree to all surrounding trees. In a first step, all trees were assessed and corrected if necessary. In a second step, these segmentations were verified and potentially refined. This was done by two authors (J. H. and J. v. D.). The corrected segmentation covers an area of 112 m by 103 m. It includes 235 trees, 156 of which are located entirely within the segment and 79 of which are partly within the segment (Figure 1b). We appended 11 m of the unlabeled forest point cloud to the labeled segment (Figure 1b, green area at the edge). This area is included in the dataset since it provides relevant context information for the manually segmented parts. The ratio of non-tree to tree points in the dataset is approximately 2:3. We make the forest plots employed in this work along with the generated segmentation labels available from https://doi.org/10.25625/VPMPID.
3 TreeLearn Pipeline
Figure 2: The pipeline for segmenting a forest point cloud. The circled numbers correspond to the five steps of the pipeline described in Section 3.

### v3

# TreeLearn: A deep learning method for segmenting individual trees from ground-based LiDAR forest point clouds

@@ Abstract
Laser-scanned point clouds of forests make it possible to extract valuable information for forest management. To consider single trees, a forest point cloud needs to be segmented into individual tree point clouds. Existing segmentation methods are usually based on hand-crafted algorithms, such as identifying trunks and growing trees from them, and face difficulties in dense forests with overlapping tree crowns. In this study, we propose TreeLearn, a deep learning-based approach for tree instance segmentation of forest point clouds. TreeLearn is trained on already segmented point clouds in a data-driven manner, making it less reliant on predefined features and algorithms. Furthermore, TreeLearn is implemented as a fully automatic pipeline and does not rely on extensive hyperparameter tuning, which makes it easy to use. Additionally, we introduce a new manually segmented benchmark forest dataset containing 156 full trees. The data is generated by mobile laser scanning and contributes to create a larger and more diverse data basis for model development and fine-grained instance segmentation evaluation. We trained TreeLearn on forest point clouds of $6665$ trees, labeled using the Lidar360 software. An evaluation on the benchmark dataset shows that TreeLearn performs as well as the algorithm used to generate its training data. Furthermore, the performance can be vastly improved by fine-tuning the model using manually annotated datasets. We evaluate TreeLearn on our benchmark dataset and the Wytham Woods dataset, outperforming the recent SegmentAnyTree, ForAINet and TLS2Trees methods. The TreeLearn code and all datasets that were created in the course of this work are made publicly available.
@@
Correspondence: jonathan.henrich@uni-goettingen.de
@@ 1 Introduction
Today, forests are not only exposed to an ever-growing pressure from human use but also to a rapidly changing environment due to climate change. To understand changes in forest composition and structure as a response to these stressors, as well as for successful forest management, scientists and forest managers require precise and detailed information on the current status of our forests. To assess the structural characteristics of forests or individual trees, laser scanning techniques are a powerful tool [1, 2, 3, 4]. They allow to create high-resolution 3D point clouds of the surfaces in a forest. To analyze such data, it is useful to separate the point cloud into the individual trees (Figure 1a), e.g. to classify them by their species [5, 6, 7] or to estimate the above-ground biomass [8, 9]. Separating a point cloud into individual trees is an instance segmentation problem: tree points must first be identified in the point cloud and then each point must be assigned to an individual tree instance. Since manually segmenting forest point clouds is highly time-consuming and subjective, there is a need for methods that automate this process. The method proposed in this work tackles this task for point clouds acquired using laser scanners that are able to capture forest structure below the canopy. Primary sources of such point clouds are terrestrial and mobile laser scanning (TLS and MLS), but also laser scanning via low-flying unmanned aerial vehicles (UAV). Point clouds obtained via airborne laser scanning (ALS) do not capture the lower parts of the forest in high detail and therefore require a different set of methods, e.g. [10, 11, 12].
@@
The automatic segmentation of MLS and TLS point clouds into individual tree instances is a relatively young research field that has so far been mostly tackled by algorithms based on hand-crafted features. These algorithms usually first identify tree trunks and then assign the remaining points to these trunks based on a fixed set of rules [13, 14]. In early works, these rules were relatively simple. For example, some authors divided the point cloud into clusters and merged them based on distance and relative orientation [15, 16]. Follow-up work building upon these methods included more elaborate local geometry and shape features [17]. More recently, Fu et al. [18] avoided complicated feature computations and proposed to cluster trees from bottom to top in a layer-by-layer manner. Another strand of literature conceptualized the forest as a graph and segmented trees via graph cut [19, 20, 21] or graph pathing methods [22]. Furthermore, some works aimed to incorporate biological theories. For example, Liu et al. [23] constructed features based on plant morphology theory which were then used in a region-growing approach. Similarly, Tao et al. [24] developed a shortest-path algorithm based on metabolic theory, which states that plants tend to minimize the transferring distance to the root. Wang [25] also employed metabolic theory by constructing a superpoint graph where the node features depend on shortest path analysis. Recently, Xu et al. [26] proposed a segmentation method that identifies distinctive features in the forest and extracts individual trees based on topological structure. Compared to previous approaches, the method is more convenient for practitioners as it reduces the need for intensive hyperparameter tuning.
@@
Although modern frameworks achieved significant improvements in performance, automated tree instance segmentation remains prone to errors [27], especially in dense forests with heavily intersecting tree crowns. In such scenarios, hand-crafted features and pre-defined heuristics appear to be too inflexible to account for the plethora of possible interactions between trees. Machine learning approaches could overcome these limitations, since features and association rules are derived directly from the data through gradient-based learning. In fact, recent advances in point cloud processing have been dominated by machine learning methods, as indicated by their performance on various point cloud processing benchmarks, e.g. [28, 29, 30].
@@
Although machine learning methods have been applied to the segmentation of trees from forest point clouds  [31, 32, 33, 34], they mostly rely on projections into 2D. As a result, they lose information, because the network does not operate on the original point cloud. Other works do employ 3D deep learning for semantic segmentation, but resort to classical algorithms for tree instance segmentation [35, 36, 37, 38]. To our knowledge, there are only a few studies in the related field of urban tree segmentation that actually employ 3D deep learning methods for tree instance segmentation [39, 40]. Furthermore, the concurrent works ForAINet [41] and SegmentAnyTree [42] recently showed the potential of directly performing deep-learning-based instance segmentation on forest point clouds. It is surprising that this research direction has not been explored more thoroughly since state-of-the-art instance segmentation methods for 3D point clouds exclusively operate on 3D input data [43, 44, 45, 46, 47, 48], making them promising candidates for improving existing tree segmentation methods. In this paper, we adapt one of these methods [44] and propose TreeLearn, a deep learning method to segment trees from forest point clouds. Our method first identifies tree points in the point cloud, then projects them towards the respective tree base, and groups them via density-based clustering. The density-based clustering is performed at once on the whole input such that complete tree instances are obtained in a single step. This is a crucial difference to the concurrent works ForAINet and SegmentAnyTree where tree instances are obtained by merging multiple incomplete tree predictions, which can lead to errors such as fragmentation. Our method is realized in an automatic pipeline which can be used out-of-the-box without extensive hyperparameter tuning. This reduces the number of work steps required to be performed by the user to a minimum, which makes our method easy to apply. Our method achieves state-of-the-art results on two tree instance segmentation datasets, outperforming the recent similar models ForAINet and SegmentAnyTree.
@@
Since data-driven deep learning methods require large amounts of data for parameter optimization, high-quality labeled forest point clouds are urgently needed. Although there are some studies that use forest point clouds with manually corrected segmentation labels, some of them do not make their datasets public [20, 21]. Furthermore, those studies that provide publicly available datasets [49, 50, 13, 11, 51] often have properties which limit their use for training machine learning methods. For example, Calders [49] only provides segmentation labels for a small number of trees located within a larger forest plot without segmentation labels. Weiser et al. [50] acknowledge quality limitations of their tree segmentations as well as low laser scanning coverage. Other studies only provide labeled trees instead of the complete point clouds [13, 52, 11]. We argue that machine learning methods should be able to handle the complete registered point clouds to make its usage as easy as possible for practitioners. Furthermore, the separation of tree points from understory and ground points is often non-trivial and therefore itself an interesting training objective. A concurrent work [51] acknowledges the need for complete labeled forest point clouds by providing a large-scale dataset with high-quality semantic and instance segmentations of 1130 trees acquired from UAV scans. In our work, we also provide a benchmark dataset of a continuous forest with high-quality manual crown segmentations and classification of points into non-tree and tree. It contains 156 trees that are completely within the labeled forest area (Figure 1b). Furthermore, we extend the publicly available labeled trees in the LAUTx [13] and Wytham Woods [52] datasets by refining the segmentation labels and propagating them to the complete point cloud. These dataset contributions are an important step towards creating a rich data basis for the development of powerful and general machine-learning-based tree segmentation algorithms. Additionally, the datasets can be used to systematically evaluate and compare existing methods.
In summary, our study has three main contributions:
@@
  • A thorough comparison between TreeLearn and other models (Lidar360, TLS2Trees, ForAINet and SegmentAnyTree) on our benchmark dataset and the Wytham Woods dataset.
2 Material and methods
In the following, we describe both the datasets used in this study and our TreeLearn approach in detail. Readers unfamiliar with the technical background can skip directly to the results section.
2.1 Datasets
The data basis for model training, validation and evaluation consists of two parts: (1) We provide MLS point clouds from 19 forest plots located in Germany (our data). For one of these plots, we performed manual segmentation, ensuring a clean and controlled separation of crowns. The other plots were automatically segmented using the Lidar360 software [53]. (2) The publically available LAUTx [13], Wytham Woods [52], and FOR-Instance [51] datasets (additional data). The LAUTx and Wytham Woods datasets do not originally contain the understory and ground component but only individual segmented trees, complicating their application for training and evaluation of our pipeline. We refined and propagated the tree labels of these datasets to the full point clouds which we obtained by contacting the authors. In Sections 2.1.1 and 2.1.2, we describe both our data and the additional data. Section 2.3 explains how these datasets were used for training, validation, and testing. All datasets that were created or extended in the course of this work are made publicly available as LAS files. LAS files allow to save point cloud coordinates together with labels. Following Puliti et al. [51], we use the treeID and classification field to distinguish tree points, non-tree points and non-annotated points.
2.1.1 Our data
Our data consists of 19 MLS point clouds located in Germany. Eight plots are located near the city of Göttingen, in Lower Saxony, two near Allstedt in Saxony-Anhalt, two near Oppershofen in Hesse, and seven near Lübeck in Schleswig-Holstein. The stands are all dominated by European beech (Fagus sylvatica L.) and between 92 and 162 years of age. The size of the study plots ranges from 1.0 to 2.2 ha. Ten plots have been managed for decades, for the others management has been abandoned for at least 30 years. More detailed information on the individual forest plots can be found in Appendix A.
Data acquisition
The plots were recorded in February 2021, in leaf-off condition, using a ZEB Horizon mobile laser scanner (Geoslam Ltd., Nottingham, UK). The device uses a laser to measure the distance and direction to surrounding objects while being carried through a scene. Based on the principle of simultaneous localization and mapping (SLAM) the ZEB Horizon creates a 3D point cloud of the scenery up to a distance of 100 m from the walking trajectory of the operator. Each hand-held walkthrough with the scanner resulted in a file that was used to create a point cloud by conducting the spatial coregistration (actual SLAM) in GeoSlam HUB Version 6 [54]. The resulting point clouds were subsampled to a resolution of 1 cm to match the effective resolution of the ZEB Horizon mobile laser scanner. Additionally, the statistical outlier removal functionality provided by CloudCompare [55] was used to remove points that are not of interest, such as scanned particles in the air.
Segmentation labels
We obtained our segmentation labels for the forest point clouds in two ways: (1) by using the automatic segmentation functionality provided by the Lidar360 software [53] and (2) by manually correcting the segmentation labels of a single forest plot.
@@ 2.1.1 Our data
To automatically segment the 19 forest point clouds into individual trees, we used the TLS-package from Lidar360 as described by Neudam et al. [56]. Lidar360 employs an implementation of the comparative shortest-path algorithm from Tao et al. [24] for segmenting the point clouds. The hyperparameters of the Lidar360 software were optimized to produce visually appealing tree segmentations. A minimum size of 10 m was chosen as a requirement for tree classification. Consequently, vegetation that is smaller than 10 m or points belonging to the forest ground were classified as non-tree points. Lidar360’s segmentation algorithm operates on a point cloud that is terrain-normalized, i. e. has a flat ground surface. This was achieved in a separate preprocessing step. We matched the terrain-normalized points with the original points based on their x- and y-coordinates to obtain labels for the unnormalized point clouds which were used in this study. The automatic segmentation contains numerous errors, including multiple trees detected as a single tree, non-tree points identified as part of trees, and inaccuracies in discerning tree boundaries in crowns and branches that are close to each other (see Section 3 for qualitative segmentation results). Since these labels contain numerous mistakes, we refer to them as noisy labels.
Manually corrected plot: L1W
@@
For one forest plot, L1W, we manually corrected the segmentation labels generated by Lidar360 with the help of CloudCompare. To do this, a python script was used to iterate through all trees in the dataset. In each iteration, the central tree was loaded together with all surrounding trees in CloudCompare and all mistakes that were associated with the central tree were corrected. This protocol ensured that mistakes were tackled one tree at a time to guide the focus of the annotators and avoid sensory overwhelm. (2) Two annotators (authors J. H. and J. v. D.) were involved in the procedure. Trees were manually corrected by one annotator and potentially refined by a second annotator. In case of ambiguities, the two annotators engaged in active discussion. The corrected segmentation covers an area of 112 m by 103 m. It includes 200 trees, 156 of which are located entirely within the segment and 44 of which are partly within the segment (Figure 1b). The partly labeled trees are also included in the dataset because they can potentially be used for model training since they contain the trunks. However, they should not be used for evaluation since parts of the crown are missing. We appended 11 m of the unlabeled forest point cloud on the edges of the labeled segment (Figure 1b, green area at the edge). This area is included in the dataset since it provides relevant context information for the manually segmented parts. The ratio of non-tree to tree points in the dataset is approximately 2:3. It should be noted that trees smaller than 10 m were assigned to the non-tree class for this dataset. This is in contrast to the other datasets used in this work (see Section 2.1.2) that also contain annotations for smaller trees.
2.1.2 Additional data: FOR-Instance, Wytham Woods and LAUTx
Table 1: Summary of the characteristics of several published labeled forest datasets. Information on NIBIO, NIBIO2, CULS, TU_WIEN, SCION and RMIT was taken from [51]. The top four rows indicate the datasets that were created or extended in the course of this work.
Name Country Reference $n$ plots $n$ trees Annotated area ($ha$) Forest type Sensor
L1W Germany Neudam et al. [56] 1 200 1.16 temperate deciduous forest GeoSLAM ZEB-Horizon
Lidar360 Germany Neudam et al. [56] 18 6665 25.09 temperate deciduous forest GeoSLAM ZEB-Horizon
LAUTx Austria Tockner et al. [13] 6 514 0.83 temperate mixed forest GeoSLAM ZEB-Horizon
Wytham England Calders et al. [52] 1 $\text{877}^{\text{a}}$ 1.52 temperate deciduous forest Riegl VZ-400
NIBIO Norway Puliti et al. [57] 20 575 1.21 coniferous dominated boreal forest Riegl miniVUX-1 UAV
NIBIO2 Norway Puliti et al. [57] 50 3062 2.87 coniferous dominated boreal forest Riegl miniVUX-1 UAV
CULS Czech Republic Kuželka et al. [58] 3 47 0.33 coniferous dominated temperate forest Riegl VUX-1 UAV
TU_WIEN Austria Wieser et al. [59] 1 150 0.55 deciduous dominated alluvial forest Riegl VUX-1 UAV
SCION New Zealand Unpublished 5 135 0.33 non-native pure coniferous temperate forest Riegl miniVUX-1 UAV
RMIT Australia Unpublished 1 223 0.37 Native dry sclerophyll eucalypt forest Riegl miniVUX-1 UAV
a The trunks of trees in Wytham Woods sometimes diverge directly at the ground. In the original publication, these parts of a tree are labeled with the same number and
0 a different character (e.g. 1a, 1b, 1c, …). Unlike other works [42], we treat these cases as separate instances so that we arrive at 877 trees.
In addition to our datasets, the existing literature was searched for further forest point clouds with tree segmentation labels. First, there is the recently published FOR-Instance dataset [51] in which tree labels and fine-grained semantic labels were manually added to point clouds from existing works. These point clouds have been captured via UAV-laser scanning and consist of diverse forest plots located in Norway (NIBIO), Czech Republic (CULS), Austria (TU_WIEN), New Zealand (SCION) and Australia (RMIT). A summary of the characteristics of each dataset can be found in Table 1. More precise information can be found in the respective publications.
Apart from these point clouds, two published datasets were identified that consist of high-quality segmented trees obtained by an automatic segmentation algorithm that were manually corrected [Wytham Woods and LAUTx; 13, 52]. The respective authors were contacted to obtain the complete unlabeled point clouds. These point clouds additionally contain non-tree points, i.e. belonging to the understory or ground, and non-annotated points, i.e. points that belong to trees but have not been annotated in the published datasets. For example, some parts of the tree crown that are hard to clearly assign to a specific tree might not have been annotated.
To obtain labels for the complete point clouds, the tree labels from the published datasets have to be propagated and the remaining points must be assigned to the classes “non-tree” or “non-annotated”. This was done as follows:
  1. 1.
    For each point in the unlabeled forest point cloud, the most common tree label within a 10 cm radius was assigned.
  2. 2.
    Among the remaining unlabeled points, non-tree points were identified using proximity-based clustering: All points that were within a 30 cm distance to each other were linked and the largest connected component was labeled as non-tree points. The large grouping radius together with the high resolution of the point clouds ensured that all understory and ground points were added to the non-tree class.
  3. 3.
    The points that were still unlabeled at this stage, i.e. not part of the largest connected component in the previous step, represent tree points that have not been annotated and were assigned to the non-annotated class. This information can be used to disregard these points during training.
  4. 4.
    Finally, we visually inspected the point clouds to ensure that they were adequately divided into trees, non-tree points and non-annotated points. Remaining errors were manually corrected within a feasible scope. Specifically, one large tree was not segmented in the original labeled data of Calders et al. [52] which we added, and the tree bases of Tockner et al. [13] were corrected since they were only roughly segmented in the original labeled data.
A visualization of the result of the above procedure is provided in Appendix A. It should be noted that in several of the third party datasets the segmentation quality decreases drastically for trees smaller than 10 m. In Wytham Woods and NIBIO2 (see Table 1) smaller trees are inconsistently labeled, i.e. sometimes as a tree and sometimes as non-tree. In LAUTx, smaller trees have severe quality limitations. For example, they are often only roughly separated from non-tree points.
2.2 TreeLearn pipeline